### v1

@@ Appendix D Extended Related Work
Recently, Steinke et al. (2023) proposed a novel scheme for auditing differential privacy with $O(1)$ training rounds. This approach enables privacy audits using multiple training examples from the same model training, if examples are included in training independently (which requires control over the training phase, and altering the target model). They demonstrate the effectiveness of this new approach on DP-SGD, in which they achieve meaningful empirical privacy lower bounds by training only one model, whereas standard methods would require training hundreds of models. Our work builds on the theory from Steinke et al. (2023), but removes the need from controlling and altering the training phase, as PANORAMIA requires no retraining.

### v2

@@ Appendix F Extended Related Work
Recently, Steinke et al. (2023) proposed a novel scheme for auditing differential privacy with $O(1)$ training rounds. This approach enables privacy audits using multiple training examples from the same model training, if examples are included in training independently (which requires control over the training phase, and altering the target model). They demonstrate the effectiveness of this new approach on DP-SGD, in which they achieve meaningful empirical privacy lower bounds by training only one model (the strongest results are achieved by including canaries in the training set though, which is not possible in our setting), whereas standard methods would require training hundreds of models. Our work builds closely on the theory from Steinke et al. (2023), but introduces a baseline model to account for distribution shifts. The key difference that enables our work to account for member/non-member distribution shifts lies in how we create the audit set in our privacy game (Definition 2). In Steinke et al. (2023), the audit set is fixed, and data points are randomly assigned to member or non-member by a Bernoulli random variable $S$. Members are actually used in training the target model $f$, while non-members are not (so assignment happens before training). In our framework, we take a set of known members (after the fact), and pair each point with a non-member(generated i.i.d. from the generator distribution). We then flip $S$ to sample which data point of each pair will be shown to the “auditor” (MIA/baseline) for testing, thereby creating the test task of our privacy measurement. When we replace our generated data with in-distribution independent non-members (PANORAMIA RM;RN), we exactly enforce the same independence as Steinke et al. (2023) (and as a result have $c=0$, and $\tilde{\epsilon}$ is a lower-bound on $\epsilon$), except that we “waste” some member data by drawing our auditing game after the fact. The difference in analysis accounts for distribution shifts when our non-members are generated.
The field of MIA without direct connections to privacy leakage measurement has also seen a lot of recent activity, with new proposals to improve the strength of MIAs in various settings (Carlini et al., 2022b; Nasr et al., 2019; Suri et al., ; Zarifzadeh et al., 2023). While it is not the focus of this paper, an interesting avenue for future work is to port ideas from these new MIAs to improve PANORAMIA’s MIA and baseline. Such transfer of MIA progress could lead to more powerful measurements with PANORAMIA. More closely related to our proposal, a recent set of works on Das et al. (2024); Duan et al. (2024); Meeus et al. (2024) MIAs on foundation models has observed that current evaluations are limited due a lack of availability of non-member data. Indeed, foundation models typical include all known data at the time of their training, and there is no well known public set of in-distribution non-member data. Evaluation tasks typical rely on more recent data, which is known to consist in non-members, but suffers from distribution shifts. Das et al. (2024); Duan et al. (2024); Meeus et al. (2024) all show that such a shift invalidates MIA evaluations by showing that “blind attacks” (i.e., our baselines) perform better than proposed MIAs. One could apply our proposed framework to such a setting: when the baseline in our framework performs better than the MIA, PANORAMIA returns a privacy loss measurement of zero instead of misleading measurements of MIA performance. We thus believe that PANORAMIA is an important step towards a theory for rigorous evaluations of membership inference for MIA models.